OpenStudy (anonymous):

Two construction companies are building 1,100 feet tall skyscrapers. Company A has finished 550 feet of construction in 5 1/2 months. Company B has finished 385 feet of construction in 3 1/2 months. If both companies continue to build at the same rate, which company will reach the top first? Pls help me😁😁😁😁

jhonyy9 (jhonyy9):

do you have same idea ?

jhonyy9 (jhonyy9):

for how begin this calculi ?

OpenStudy (anonymous):

Can you explain how to solve this problem and an explanation

jhonyy9 (jhonyy9):

first you need calculi for company A and B in how many day building 1 feet can you calculi it here now ?

jhonyy9 (jhonyy9):

5 1/2 months = ? days

OpenStudy (anonymous):

31 times 5 1/2 Right?????

OpenStudy (anonymous):

Or 30 times 5 1/2

jhonyy9 (jhonyy9):

suppose yes than you get one month has 30 or 31 days - depend how you calculi

OpenStudy (anonymous):

Ok thanks I think I got it

jhonyy9 (jhonyy9):

ok. but i like check your calculi here - ok. ?

jhonyy9 (jhonyy9):

and again than you see this 550 feet is a half part of this needed building feet ,yes ?

jhonyy9 (jhonyy9):

so this mean that than you duble this period of 5 1/2 month you get the necessary tome for company A to building these 1100 feet - yes ?

jhonyy9 (jhonyy9):

sorry - the necessary time ...

jhonyy9 (jhonyy9):

ok. so than doubling this period of company A you get 1100 feet ended in 11 month - yes ?

jhonyy9 (jhonyy9):

so for company B you know that build 385 feet in 3 1/2 month so 3 1/2 month are 105 days 385 feet in 105 days 1100 feet --- x days ---------------------- cross multiplie and calculi x x = 1100*105/385 = 300 days - this mean that company B build these 1100 feet in 300 days

jhonyy9 (jhonyy9):

for company A this 11 month mean 11*30=330 days so from this result that company B will reach the top first ok ? hope you understand it sure yes ?

OpenStudy (mathmale):

This problem invovles rates of change: one constant rate for one company and another, different, constant rate for the other company. We assume that at the beginning of construction, the height of the building is 0, and at the end the height is 1,100 feet. An appropriate model for this construction would be a linear function with zero initial value (height of building is 0 ft. and time t=0). Building 1: \[h _{A}(t)=\frac{550ft }{ 5.5 mo }*t _{A}=1100ft\]

OpenStudy (mathmale):

Constructing a similar equation for Building B:\[h _{B}(t)=\frac{385ft }{ 3.5 mo }*t _{B}=1100ft\]

OpenStudy (mathmale):

Solve each of these two equations for the time required to complete all 1100 feet of construction. The contractor with the shorter construction time "wins."

OpenStudy (mathmale):

Note how these equations follow the pattern Amount = (rate)*(time). Also note that you could solve both equations for time before inserting numeric values: time = (amount) / (rate)

OpenStudy (mathmale):

Note that there's no need to convert "months" to "days."